# ProDAG: Projected Variational Inference for Directed Acyclic Graphs

### v5

@@ 1 Introduction
$\tilde{W}\sim\mathbb{P},\quad\lambda\sim\pi,\quad W=\operatorname{pro}_{\lambda}(\tilde{W}),$ (1)
@@
Figure 1: Illustration of ProDAG’s projected distributions. Samples (blue dots) from an unconstrained continuous space (blue ellipse) are projected onto the nearest acyclic matrix within an $\ell_{1}$-constrained region (orange diamonds). Projected samples (orange dots) satisfy acyclicity and sparsity constraints. Theoretically, we show that for any continuous $\tilde{W}\sim\mathbb{P}$ the function $\operatorname{pro}_{\lambda}(\tilde{W})$ is uniquely defined and measurable. This implies that the projected distribution is a valid distribution over DAGs.
@@ 2 Related work
As mentioned, existing approaches for Bayesian inference on DAGs broadly fall into two groups: (1) regularization-based methods that include a continuous acyclicity penalty via a Gibbs-type prior in the variational objective (Annadani et al., 2021; Lorch et al., 2021; Geffner et al., 2022), and (2) state-augmentation methods that, e.g., place a prior over the graph’s topological ordering (Cundy et al., 2021; Charpentier et al., 2022; Bonilla et al., 2024; Toth et al., 2024). The node-potential augmentation of Annadani et al. (2023a) is related to the latter category. See also Wang et al. (2022); Deleu et al. (2023) for approaches based on probabilistic circuits and generative flow networks. Our approach differs by designing a prior and approximate posterior directly over DAGs without modeling the (discrete) topological ordering and without using a penalty that does not place mass on zeros.
@@ 3.1 Description
To recap the introduction, we induce a distribution over weighted acyclic adjacency matrices $W$ as $W=\operatorname{pro}_{\lambda}(\tilde{W})$, where $\operatorname{pro}_{\lambda}$ is defined in (2), $\tilde{W}\sim\mathbb{P}$ is a continuous matrix, and $\lambda\sim\pi$ bounds the $\ell_{1}$ norm of $W$ [^2] The distribution $\pi$ can be a Dirac for a distribution over graphs of fixed sparsity (same number of edges) or an exponential for a distribution over graphs of varying sparsity. Since acyclic graphs have no self-loops (i.e., $W$ always has a zero diagonal), we can take $\mathbb{P}$ to be any continuous distribution over the off-diagonal elements of $\tilde{W}$ (e.g., a multivariate Gaussian), with the diagonal elements fixed as zero. The matrix $\tilde{W}$ can thus be interpreted as the weighted adjacency matrix of a simple directed graph, with the projection removing all cycles.
@@
Conceptual interpretation
The resulting distribution $p(W)$ can be understood as a mix of continuous components and exact zeros. To elaborate, first observe that the nonzeros in $W$ may be indexed by a sparsity pattern $S\in\{0,1\}^{p\times p}$ (i.e., an unweighted acyclic adjacency matrix). The nonzeros are then $W_{S}$, where $W_{S}$ is the restriction of $W$ to $S$. Because the $\ell_{1}$ constraint shrinks the nonzeros towards zero, the continuous $W_{S}$ is a transformation $f_{\lambda}$ of $\tilde{W}_{S}$, i.e., $W_{S}=f_{\lambda}(\tilde{W}_{S})$, where $f_{\lambda}$ is implicit in the projection. This mixture interpretation suggests an interesting representation of $p(W)$:
@@ 3.2 Properties
Theorem 1 states that, for any continuous $\tilde{W}\sim\mathbb{P}$, the function $\operatorname{pro}_{\lambda}(\tilde{W})$ is uniquely defined and measurable. This important result implies that a projected distribution is a valid distribution since a measurable function $g$ applied to a random variable $z$ produces a well-defined random variable $g(z)$.
@@ 4.3 Gradients
In many Bayesian learning frameworks, including our variational framework, it is necessary to compute gradients. In our case, this computation is complicated by the presence of the projection since we need its gradients with respect to $\tilde{W}$ and $\lambda$. Though it is technically possible to differentiate through our algorithms using automatic differentiation, it is practically infeasible due to the iterative nature of the path-following algorithm. As it turns out, it is possible to analytically evaluate the gradients via the implicit function theorem. This technique involves differentiating the Karush-Kuhn-Tucker (KKT) conditions of the projection that implicitly define $W$ as a function of $\tilde{W}$ and $\lambda$. The gradients from this implicit differentiation technique are in Appendix C alongside their derivation.
@@ 5.1 Prior and variational posterior
Let $p(\tilde{W},W\mid X)$ be the true joint posterior and let $q_{\theta}(\tilde{W},W)$ be the variational joint posterior. Since $W$ is a deterministic function of $\tilde{W}$, the variational joint posterior is the product of the variational marginal posterior $q_{\theta}(\tilde{W})$ and a Dirac delta function: $q_{\theta}(\tilde{W},W)=q_{\theta}(\tilde{W})\delta(W-\operatorname{pro}_{\lambda}(\tilde{W}))$. The marginal $q_{\theta}(\tilde{W})$ can be a continuous distribution such as a multivariate Gaussian. In that case, the parameters $\theta$ contain means and covariances. Choosing $q_{\theta}(\tilde{W})$ is the same as choosing $\mathbb{P}$ in (1). For simplicity and clarity of exposition, we treat the sparsity parameter $\lambda$ as fixed, i.e., $\pi$ in (1) is a Dirac.
@@ 5.2 Evidence lower bound
To estimate this expectation, we sample DAGs from $q_{\theta}(W)$ and evaluate the log-likelihood on these samples. To sample from $q_{\theta}(W)$, we need only sample from $q_{\theta}(\tilde{W},W)$ and discard $\tilde{W}$. The second term on the right-hand side of (5) is the KL divergence between the variational posterior $q_{\theta}(\tilde{W},W)$ and the prior $p(\tilde{W},W)=p(\tilde{W})\delta(W-\operatorname{pro}_{\lambda}(\tilde{W}))$. Since $W$ is a deterministic function of $\tilde{W}$, it is not too difficult to see that the KL divergence between the joint posterior and prior is equal to the KL divergence between the marginal posterior and prior on $\tilde{W}$. Hence, the second term also simplifies as
$\operatorname{KL}[q_{\theta}(\tilde{W},W)\parallel p(\tilde{W},W)]=\operatorname{KL}[q_{\theta}(\tilde{W})\parallel p(\tilde{W})].$
Provided that $q_{\theta}(\tilde{W})$ and $p(\tilde{W})$ are chosen from distributional families with closed-form density functions such as multivariate Gaussian distributions, this term is amenable to analytical evaluation. The gradients of Gaussians and other (more complex) distributions for $\tilde{W}$ are readily accommodated in the ELBO using the well-known reparameterization trick (see, e.g., Kingma and Welling, 2014).
@@ 5.3 Algorithm
With above remarks in mind, we maximize the ELBO using the gradient descent-based Algorithm 1.
@@
     Sample $\tilde{W}^{(l)}\sim q_{\theta}(\tilde{W})$ and set $W^{(l)}=\operatorname{pro}\lambda(\tilde{W}^{(l)})$ for $l=1,\ldots,L$
     Compute $\hat{\operatorname{ELBO}}(\theta)=1/L\sum_{l=1}^{L}\log p(X\mid W^{(l)})-\operatorname{KL}[q_{\theta}(\tilde{W})\parallel p(\tilde{W})]$
@@ 6.2 Posterior learning
$\tilde{\omega}\sim\mathbb{P},\quad\lambda\sim\pi,\quad\omega=\operatorname{pro}_{\lambda}(\tilde{\omega}),$
@@
$\operatorname{ELBO}(\theta)=\operatorname{E}_{q_{\theta}(\omega)}[\log p(X\mid\omega)]-\operatorname{KL}[q_{\theta}(\tilde{\omega})\parallel p(\tilde{\omega})],$
@@ 7.3 Nonlinear synthetic data
Figure 5 reports results on datasets simulated from nonlinear DAGs.
@@
Following the setup in Zheng et al. (2020), each variable in the graph is generated from a single hidden-layer neural network whose inputs are that variable’s parents. The learning task is more challenging here than in the linear setting (roughly 10× more unknown parameters for a graph of the same size), so we study datasets simulated from graphs with $p=10$ nodes and $s=20$ edges. BOSS cannot model nonlinear DAGs, so it is excluded. For the most part, the story here is similar to the linear setting: ProDAG delivers strong performance and improves over the frequentist and Bayesian baselines most of the time. Again, it particularly excels at uncertainty quantification. The results for DiBS approximately match those in Lorch et al. (2021), where its expected SHD is similar to that of the null graph. Annadani et al. (2023a) studied BayesDAG on a sample size of $n=5000$, much bigger than the largest sample size here. ProDAG does not require a sample of similarly large size to produce high-quality inference.
@@ 7.5 Additional experiments
Appendix G reports more results from experiments on additional types of synthetic data. These experiments cover (i) denser graphs, (ii) non-Gaussian noise distributions, and (iii) scale-free graphs.
@@ 8 Concluding remarks
DAGs are complex structures due to their acyclic nature, making Bayesian inference a formidable task. ProDAG is a novel approach that employs new projected distributions that have support directly on the space of DAGs. We study key properties of these distributions and show that they facilitate state-of-the-art Bayesian inference for linear and nonlinear DAGs in a scalable variational framework.
@@ Appendix A Proof of Theorem 1
$\min_{\begin{subarray}{c}S\in\{0,1\}^{p\times p},\,S\in\mathrm{DAG}\\ V\in\mathbb{R}^{p\times p},\,\|V\|_{\ell_{1}}\leq\lambda\end{subarray}}\,f_{S}(V;\tilde{W})\Longleftrightarrow\min_{\begin{subarray}{c}S\in\{0,1\}^{p\times p}\\ S\in\mathrm{DAG}\end{subarray}}\,\min_{\begin{subarray}{c}V\in\mathbb{R}^{p\times p}\\ \|V\|_{\ell_{1}}\leq\lambda\end{subarray}}\,f_{S}(V;\tilde{W}).$
Consider the inner optimization problem on the right-hand side. For any $\tilde{W}$, let $V^{\star}(\tilde{W})$ denote an optimal solution to this problem, and define the set of $\tilde{W}$ with non-unique solutions as
$A_{S}:=\{\tilde{W}\in\mathbb{R}^{p\times p}\mid f_{S}(V_{1}^{\star}(\tilde{W});\tilde{W})=f_{S}(V_{2}^{\star}(\tilde{W});\tilde{W}),\,V_{1}^{\star}(\tilde{W})\neq V_{2}^{\star}(\tilde{W})\}.$
Since $f_{S}(V;\tilde{W})$ is strictly convex in $V$ and the $\ell_{1}$ ball is convex, the optimal solution $V^{\star}(\tilde{W})$ is unique for all $\tilde{W}$. Hence, the set $A_{S}$ is empty. Now, define the set of $\tilde{W}$ with non-unique solutions across any two different adjacency matrices $S_{1}$ and $S_{2}$ as
$A_{S_{1},S_{2}}:=\{\tilde{W}\in\mathbb{R}^{p\times p}\mid f_{S_{1}}(V_{1}^{\star}(\tilde{W});\tilde{W})=f_{S_{2}}(V_{2}^{\star}(\tilde{W});\tilde{W}),\,V_{1}^{\star}(\tilde{W})\neq V_{2}^{\star}(\tilde{W})\}.$
Because $V_{1}^{\star}(\tilde{W})\neq V_{2}^{\star}(\tilde{W})$, there is at least one coordinate $(j,k)$ on which they differ. Equality of the two quadratic objectives then forces $\tilde{W}$ to lie on a single affine hyperplane in $\mathbb{R}^{p\times p}$. Such a hyperplane has measure zero when $\tilde{W}$ is endowed with a continuous probability measure. Moreover, because there are finitely many pairs $(S_{1},S_{2})$ with $S_{1}\neq S_{2}$, the union of $A_{S_{1},S_{2}}$ over all pairs also has measure zero. We thus conclude that projection (2) is unique almost surely with respect to $\tilde{W}$.
@@
$\operatorname{pro}_{\lambda}(\tilde{W})=g_{\lambda}(\tilde{W},h_{\lambda}(\tilde{W})).$
We now demonstrate that each of these functions are measurable in $\tilde{W}$. First, for fixed $S$, observe that the function $g_{\lambda}(\tilde{W},S)$ is the minimizing argument of a strictly convex optimization problem (with a compact feasible set) and $\tilde{W}\mapsto 1/2\|\tilde{W}-S\odot V\|_{F}^{2}$ is continuous in $\tilde{W}$. Hence, by Berge’s maximum theorem, $g_{\lambda}(\tilde{W},S)$ is continuous in $\tilde{W}$. Continuous functions are measurable with respect to their argument, and therefore $g_{\lambda}(\tilde{W},S)$ is measurable in $\tilde{W}$. Second, since the set of binary adjacency matrices is finite, the function $h_{\lambda}(\tilde{W})$ takes the pointwise minimum of finitely many measurable functions and returns the adjacency matrix indexed by that minimum (breaking any ties per the lexicographic order). A pointwise minimum of measurable functions is itself measurable. Lastly, since $g_{\lambda}(\tilde{W},S)$ and $h_{\lambda}(\tilde{W})$ are both measurable in $\tilde{W}$, their composition $g_{\lambda}(\tilde{W},h_{\lambda}(\tilde{W}))$ must also be measurable in $\tilde{W}$. We thus conclude that projection (2) is measurable with respect to $\tilde{W}$.

### v6

@@ 1 Introduction
[…]
@@
[…]
@@ 2 Related work
Existing approaches for Bayesian inference on DAGs broadly fall into two groups: (1) regularization-based methods that include a continuous acyclicity penalty via a Gibbs-type prior in the variational objective (Annadani et al., 2021; Lorch et al., 2021; Geffner et al., 2022), and (2) state-augmentation methods that, e.g., place a prior over the graph’s topological ordering (Viinikka et al., 2020; Cundy et al., 2021; Charpentier et al., 2022; Bonilla et al., 2024; Toth et al., 2024). The node-potential augmentation of Annadani et al. (2023a) is related to the latter category. See also Wang et al. (2022) and Deleu et al. (2023) for approaches based on probabilistic circuits and generative flow networks. Our approach differs by designing a prior and approximate posterior directly over DAGs without modeling the (discrete) topological ordering or using a penalty that does not place mass on zeros.
@@ 3.1 Description
[…]
@@
On a conceptual level, the resulting distribution $p(W)$ can be understood as a mix of continuous components and exact zeros. [^3] This mixture means that the projected distribution is multimodal, as different regions of the continuous distribution map to distinct acyclic graphs (except in degenerate point-mass cases). To elaborate, first observe that the nonzeros in $W$ may be indexed by a sparsity pattern $S\in\{0,1\}^{p\times p}$ (i.e., an unweighted acyclic adjacency matrix). The nonzeros are then $W_{S}$, where $W_{S}$ is the restriction of $W$ to $S$. Because the $\ell_{1}$ constraint shrinks the nonzeros towards zero, the continuous $W_{S}$ is a transformation $f_{\lambda}$ of $\tilde{W}_{S}$, i.e., $W_{S}=f_{\lambda}(\tilde{W}_{S})$, where $f_{\lambda}$ is implicit in the projection. This mixture interpretation suggests an interesting representation of $p(W)$:
@@ 3.2 Properties
[…]
@@ 4.3 Gradients
In many Bayesian learning frameworks, including our variational framework, it is necessary to compute gradients. In our case, this computation is complicated by the presence of the projection since we need its gradients with respect to $\tilde{W}$ and $\lambda$. Though it is technically possible to differentiate through our algorithms using automatic differentiation, it is practically infeasible due to the iterative nature of the path-following algorithm. As it turns out, it is possible to analytically evaluate the gradients via the implicit function theorem. This technique involves differentiating the Karush-Kuhn-Tucker (KKT) conditions of the projection that implicitly define $W$ as a function of $\tilde{W}$ and $\lambda$. Intuitively, the KKT conditions express how optimality of $W$ is preserved under small perturbations of $\tilde{W}$ or $\lambda$, and the implicit function theorem formalizes this dependence to yield local expressions for how $W$ varies. The gradients from this technique are in Appendix C alongside their derivation.
@@ 5.1 Prior and variational posterior
[…]
@@ 5.2 Evidence lower bound
To estimate this expectation, we sample DAGs from $q_{\theta}(W)$ and evaluate the log-likelihood on these samples. To sample from $q_{\theta}(W)$, we need only sample from $q_{\theta}(\tilde{W},W)$ and discard $\tilde{W}$. The second term on the right-hand side of (5) is the KL divergence between the variational posterior $q_{\theta}(\tilde{W},W)$ and the prior $)p(\tilde{W},W)=p(\tilde{W})\delta(W-\operatorname{pro}_{\lambda}(\tilde{W}$)). Since $W$ is a deterministic function of $\tilde{W}$, it is not too difficult to see that the KL divergence between the joint posterior and prior is equal to the KL divergence between the marginal posterior and prior on $\tilde{W}$. Hence, the second term also simplifies as
[…]
@@ 5.3 Algorithm
In light of the above remarks, we maximize the ELBO using the gradient descent-based Algorithm 1.
@@
[…]
@@ 6.2 Posterior learning
[…]
@@
[…]
@@ 7.3 Nonlinear synthetic data
Figures 5 and 6 report results on datasets simulated from nonlinear DAGs.
@@
Figure 6: Performance on synthetic datasets generated from nonlinear Erdős–Rényi DAGs with $p=20$ nodes, $s=40$ edges, and Gaussian noise. The averages (solid points) and standard errors (error bars) are measured over 10 independently and identically generated datasets.
Following the setup in Zheng et al. (2020), each variable in the graph is generated from a single hidden-layer neural network whose inputs are that variable’s parents. The learning task is more challenging than in the linear setting (more than 10× parameters in each graph), so we study graphs with $p=10$ and $p=20$ nodes ($s=20$ and $s=40$ edges). BOSS cannot model nonlinear DAGs, so it is excluded. DiBS and DiBS+ exceed available memory when $p=20$, so are excluded in that setting. ProDAG delivers strong performance and improves over the frequentist and Bayesian baselines most of the time. Again, it particularly excels at uncertainty quantification. The results for DiBS approximately match those in Lorch et al. (2021), where its expected SHD is similar to that of the null graph. Annadani et al. (2023a) studied BayesDAG on a sample size of $n=5000$, much bigger than the largest sample size here. ProDAG does not require a sample of similarly large size to produce high-quality inference.
@@ 7.5 Additional experiments
A suite of additional experiments are included in Appendix G. In particular, Appendix G.1 presents results from further synthetic experiments covering (i) denser graphs, (ii) scale-free graphs, (iii) non-Gaussian noise distributions, and (iv) heteroscedastic noise distributions. Appendix G.2 reports results on semi-synthetic datasets generated from a large real-world medical diagnostic network. Appendix G.3 provides comparisons with a Markov chain Monte Carlo (MCMC) baseline. Finally, Appendix G.4 compares the run times of ProDAG with those of its primary Bayesian competitors.
@@ 8 Concluding remarks
DAGs are complex structures due to their acyclic nature, making Bayesian inference a formidable task. ProDAG is a novel approach that employs new projected distributions that have support directly on the space of DAGs. We study key properties of these distributions and show that they facilitate state-of-the-art Bayesian inference for linear and nonlinear DAGs in a scalable variational framework. While ProDAG is computationally costlier than frequentist methods, it is well-suited to settings where reliable posterior uncertainty is critical (e.g., decision making), where limited data make sparsity valuable (e.g., biology), or where strict acyclicity and interpretability are essential (e.g., policy).
As with other approaches to structure discovery using purely observational data, identifiability of the DAG structure in ProDAG depends on the correspondence between the assumed likelihood and the true data generating process. When the likelihood correctly specifies the noise distribution implied by the underlying SEM, standard identifiability results apply under assumptions such as causal sufficiency, the Causal Markov condition, and faithfulness. For instance, in a linear SEM with homoscedastic Gaussian noise, the DAG becomes fully identifiable (Peters and Bühlmann, 2014). In this setting, which our experiments cover, ProDAG’s posterior increasingly concentrates on the correct DAG as the sample size grows. Conversely, if the likelihood is misspecified, different graphs may induce indistinguishable likelihoods, leading to a loss of identifiability. In such cases, ProDAG represents the remaining structural ambiguity through the spread of its posterior distribution.
Recent work has extended continuous acyclicity approaches to interventional data (Brouillard et al., 2020; Faria et al., 2022; Lopez et al., 2022). Though our focus is observational data, ProDAG could be adapted for interventional settings by replacing the observational likelihood with one appropriate for interventions while leaving the projected distributions unchanged. Developing this extension and assessing ProDAG in interventional contexts offers a natural and important next step for research.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately reflect the remainder of the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitations are discussed in Appendix H.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: Assumptions where relevant are stated. Proofs are provided in the appendices.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Core details for reproducing the experimental results are included in Appendices E and F. Code for exactly reproducing the experimental results is released.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Code, data, and instructions for reproducing the experiments is available.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Experimental settings are detailed in Appendix E.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: One standard errors of the mean are reported for all experiments.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Computational resources are described in Appendix E.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: The paper conforms with the NeurIPS Code of Ethics.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: Societal impacts are discussed in Appendix I.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: The paper poses no such risks.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All existing assets are cited in the paper. Version numbers and URLs are provided. The license name is included for each asset. See Appendix E.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: A package ProDAG is released with accompanying documentation.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: The paper does not involve crowdsourcing.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: The paper does not involve research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: The core method development does not involve LLMs.
@@ Appendix A Proof of Theorem 1
[…]
@@
$).\operatorname{pro}_{\lambda}(\tilde{W})=g_{\lambda}(\tilde{W},h_{\lambda}(\tilde{W}$)).
[…]